# An Investigation of FP8 Across Accelerators for LLM Inference

### v2

@@ 5.4 Decode phase
Table 4: Thin GEMM throughputs in TFLOPS for the Gaudi 2 and H100 measuring $(M\times K)\times(K\times N)$ GEMMs, where $M$ corresponds to typical batch sizes encountered during inference and $K,N$ represent hidden dimension sizes. Row-wise scaling is used for FP8 GEMMs. Throughput scales linearly with $M$ on both devices, but the Gaudi 2 consistently outperforms the H100, even for BF16. These results highlight the superior efficiency of the Gaudi 2 for thin GEMMs, crucial for LLM decoding.
@@
Figure 5: Thin GEMM MFU comparison between the Gaudi 2 and H100 for both BF16 and FP8. The Gaudi 2 maintains a similar MFU for BF16 and FP8 of similar shapes while there is a noticeable drop for the H100. The MFU differences at the same shapes are enough to provide superior TFLOPS for the Gaudi 2 over the H100 as shown in Table 4.
6 Conclusion
In this study, we provide a comprehensive analysis of FP8, an emerging numerical format that remains underexplored. We examine various FP8 scaling factor configurations and analyze LLM inference workloads based on the constraints of real-world data center environments. We make a clear exposition of the key constraints in LLM inference, also providing a framework to make precise calculations for inference throughput. Our analysis of the two phases of LLM inference highlights the strong dependence of LLM inference on hardware characteristics.
However, beyond hardware development, software optimizations that fully leverage accelerator capabilities are equally essential to maximizing performance. One key limitation of the Gaudi is its limited software support and integration. For instance, many kernel fusion techniques for dynamic FP8 scaling are currently unavailable on the Gaudi. Additionally, the decode throughputs reported in our study relied on static memory allocation as the paged attention implementations remained uncompetitive despite recent optimizations (Lee et al., 2024). Even for static memory allocation, the observed attention throughputs are lower than what would be expected from the hardware specifications. Furthermore, while not a primary focus of this work, the lack of KV cache quantization support present challenges for deployment.
Despite these concerns, the growing adoption of inference-time scaling via chain-of-thought (CoT) (Wei et al., 2022) prompting suggests that LLM workloads will increasingly be dominated by decoding. Emerging techniques that decouple the prefill and decode phases (Patel et al., 2024; Zhong et al., 2024) enable hardware to be allocated to the tasks for which they are best suited. Given our findings that the Gaudi 2 demonstrates superior decoding efficiency for FP8, this has potential implications for cloud service providers seeking cost-effective inference solutions.
Appendix A Appendix
Table 5: MMLU result for per-row, per-tensor options
  model Rowwise Scaling accuracy
Activation Weight
Llama-3.2-1B 35.12%
34.69%
34.36%
34.65%
Llama-3.2-3B 53.90%
53.92%
53.83%
53.75%
Llama-3.1-8B 62.95%
62.92%
62.80%
62.90%
Llama-3.1-70B 75.25%
75.23%
75.32%
75.39%
Table 6: MMLU 5-shot CoT for stochastic rounding using E4M3.
  Model Stochastic Rounding MMLU accuracy
Activation Weight
Llama-3.2-1B 33.94%
34.31%
35.20%
34.65%
Llama-3.2-3B 54.09%
53.95%
54.41%
53.75%
Llama-3.1-8B 62.24%
62.43%
62.55%
62.90%
Llama-3.1-70B 75.05%
75.18%
75.30%
75.39%
Table 7: Throughput and Power Measurement of Llama Instruct Models
  Model Name Mode Batch Size Seq Len Power(W) Prefill TFLOPS/HPU
Llama-3.2-1B-Instruct FP8-Static 4 512 509 412
Llama-3.2-1B-Instruct FP8-Static 4 1024 509 411
Llama-3.2-3B-Instruct FP8-Static 4 512 498 518
Llama-3.2-3B-Instruct FP8-Static 4 1024 511 537
Llama-3.1-8B-Instruct FP8-Static 4 512 471 588
Llama-3.1-8B-Instruct FP8-Static 4 1024 436 595
Table 8: Gaudi 2 throughput in TFLOPS for scaled FP8 GEMM for square matrices of the given size, excluding quantization overhead. Measured throughput relative to the peak FP8 throughput (865 TFLOPS) is included in parentheses. Hardware acceleration is only available for per-tensor scaling.
Type Size Per-Row Per-Tensor HW Accel.
E4M3 1K 494 (57.1%) 494 (57.1%) 494 (57.1%)
2K 506 (58.5%) 641 (74.1%) 641 (74.2%)
4K 735 (84.9%) 796 (92.1%) 801 (92.6%)
8K 742 (85.7%) 822 (95.0%) 852 (98.4%)
E5M2 1K 306 (35.4%) 494 (57.1%) 493 (57.0%)
2K 506 (58.5%) 642 (74.2%) 642 (74.2%)
4K 735 (84.9%) 802 (92.7%) 802 (92.7%)
8K 726 (83.9%) 825 (95.4%) 825 (95.4%)
Table 9: H100 throughput in TFLOPS for scaled FP8 GEMM for square matrices of the given size, excluding quantization overhead. Measured throughput as a ratio of peak FP8 throughput (1989.9 TFLOPS) is included in parentheses. We only include results for E4M3 because we were unable to find libraries exposing low-level E5M2 GEMM APIs on NVIDIA GPUs.
Accum. Size Per-Row Per-Tensor
FP32 1K 217 (11.0%) 186   (9.4%)
2K 299 (15.1%) 840 (42.4%)
4K 362 (18.3%) 1099 (55.5%)
8K 396 (20.0%) 1300 (65.7%)
Fast 1K 237 (12.0%) 147   (7.4%)
2K 810 (40.9%) 896 (45.3%)
4K 1136 (57.4%) 1205 (60.9%)
8K 1123 (56.8%) 1388 (70.1%)

### v3

@@ 5.6. Thin GEMM with FP8 quantization
Table 6. Thin GEMM throughputs (in TFLOPS) for the Gaudi 2 and H100 measuring $(M\times K)\times(K\times N)$ GEMMs, where $M$ corresponds to typical batch sizes during inference and $K,N$ represent hidden dimension sizes. Row-wise scaling is used for FP8 GEMMs. Throughput scales linearly with $M$ on both devices, but the Gaudi 2 consistently outperforms the H100, even for BF16. Results highlight the superior efficiency of the Gaudi 2 for thin GEMMs, crucial for LLM decoding.
@@
These differences in FP8 thin GEMM performance are likely attributed to the architectural differences of the hardware platforms. In the case of NVIDIA Hopper, details of the Tensor Core architecture remain undisclosed, making it difficult to analyze the actual implementation. We expect that future disclosure of internal implementation details would enable a more systematic investigation. In contrast, Gaudi2 provides public documentation that contains their Matrix Multiplication Engine (MME) architecture, which is an output-stationary systolic array.
Figure 7 depicts the differences between the MME architectures of the Gaudi2 and those of other commercial hardware solutions. Unlike other commercial hardware offerings that employ multiple small compute units, Gaudi2 integrates a single large MME. In architectures that rely on several small units, each unit consumes input data in every cycle, incurring significant overhead in unit-to-unit data movement. In contrast, one large MME consumes a much smaller amount of input data per cycle and requires less expensive ALU-to-ALU communication within a whole MME. The effect of communication cost leads to notable MFU differences, as shown in Figure 6. The data movement latency is effectively hidden in compute-bound conditions, but it becomes critical in memory-bound thin GEMM operations, where communication overhead can dominate performance and markedly degrade overall performance.
Furthermore, Figure 8 illustrates the reconfigurable MME sizes of Gaudi 2. While the Gaudi2 architecture contains two MMEs with size $256\times 256$, they can be folded and integrated in different sizes with a smallest width 128. The size of the MME is optimized at run time via the Gaudi Graph Compiler. This flexibility allows the hardware to adapt to a wide range of GEMM dimensions, offering substantial potential for improving LLM inference efficiency.
5.7. Exponential function optimization
While the Gaudi 2 outperforms the H100 in the LLM decode phase for short sequence lengths, its throughput gains diminish at longer sequence lengths. We find that this decline is primarily due to the absence of the dedicated Special Function Units (SFUs, also referred to as multi-function units, or MUFUs) in the Gaudi 2 & 3 to accelerate the exponential calculations in the softmax.
During typical LLM inference, SFUs perform crucial tasks. Although GEMM operations that leverage tensor cores dominate LLM workloads, specialized functions such as softmax and SiLU are also necessary. For decoding workloads, the ratio of exponential computations relative to GEMM computations increase in proportion to the sequence length. As a result, the softmax operation becomes a bottleneck for decoding at long sequence lengths. GPUs equipped with SFUs can process exponential functions on SFUs while simultaneously executing GEMMs on tensor cores, effectively parallelizing the workload and boosting overall performance.
In contrast, both the Gaudi 2 & 3 have a disadvantage in that they lack SFUs and must handle exponential computations via their TPC cores. They also have a lower ratio of vector core compute relative to matrix computation units compared to the H100. Peak TPC core BF16 throughput is 11 TFLOPS for the Gaudi 2 and 28.7 TFLOPS for the Gaudi 3, whereas the H100 has 133.8 TFLOPS BF16 CUDA core throughput and includes separate SFU cores. This implies that despite the notable efficiency of thin GEMMs in the Gaudi 2, exponential function calculations emerge as the main bottleneck for long sequences, as the softmax computation scales $O(BS)$ for batch size $B$ and sequence length $S$ during decoding.
The Gaudi 2 incorporates an embedded look-up table (LUT), which could theoretically store precomputed special-function values via piecewise linear approximation. If the LUT offers sufficiently low access latency and can hold a large enough range of values, it may be a valid substitute for SFUs. However, Gaudi 2 currently lacks software support for programming the LUT. Although it has strong potential to improve performance with long sequences, the limited software stack constrains its practical effectiveness.
Figure 7. Gaudi2 employs one large systolic array unit for each MME, unlike the many other commercial hardware employs many small units.
Figure 8. Dynamically reconfigurable MME architecture on Gaudi for target GEMM optimization.
6. Throughput shifts the TCO balance
As described in Section 2, the TCO varies significantly depending on factors such as data precision, workload characteristics, and batch size. Figure 9 illustrates how the TCO ratio changes under varying conditions, particularly focusing on two key parameters: the throughput ratio and the server cost ratio. In our TCO model, system A represents the Gaudi 2 and system B represents the H100. In Figure 9, movement along the x-axis reflects the server cost ratio, which is determined by factors such as device procurement cost, infrastructure constraints, and other datacenter-specific considerations. This ratio varies across CSPs depending on their individual deployment environments. Therefore, the dominant factor influencing the TCO ratio is the throughput ratio ($R_{Th}$).
In the LLM decoding phase, which dominates overall inference latency, the Gaudi 2 achieves nearly 150% the throughput of BF16 when using FP8 quantization, while the H100 shows little to no performance gain. As a result, $R_{Th}$ increases, and the TCO ratio shifts upward along the y-axis, approaching the green region in the figure, indicating higher cost efficiency for the Gaudi 2. Conversely, in scenarios involving long sequences, the Gaudi 2 suffers from the softmax bottleneck due to its lack of a dedicated SFU or an alternative workaround. This lowers the throughput ratio, shifting the TCO ratio downward along the y-axis toward the red region, which implies reduced cost efficiency for the Gaudi 2.
Ultimately, the actual position on the TCO map depends on the specific conditions of each datacenter and on the specific operations required by the CSP. Therefore, accurate evaluation requires a detailed analysis of the operational environment and workload profile unique to each CSP.
Figure 9. Visualization of TCO Variation by Server Cost and Throughput Ratios.
7. Conclusion
In this study, we present a scalable framework for comparing AI accelerator options in terms of datacenter TCO. Using our model, we examine two key factors that influence throughput and are directly connected to the TCO: numerical precision and LLM inference workload characteristics. We evaluate two representative next-generation accelerators, the NVIDIA H100 and Intel Gaudi 2, and analyze their behavior under FP8 quantization.
However, our study has several limitations. First, our analysis is based on a case study focused specifically on Llama-based models in the context of inference. As such, the results may not directly generalize to other workloads, such as training. Second, we adopt FP8 precision to reflect anticipated trends in future-generation models. However, the applicability of our findings depends on the extent of hardware support for FP8, which remains a limitation.
While our study provides valuable insights into cost-performance trade-offs, additional research is required to generalize the results to full-scale datacenter environments with diverse infrastructure, workload mixes, and deployment strategies. We hope that the insights provided in this study help inform deployment decisions and guide future accelerator design efforts aimed at improving the TCO of LLM inference in datacenter environments.